OpenStudy (anonymous):

sqrt12a^5b*sqrt8a^6b^4

OpenStudy (saifoo.khan):

whole thing?

OpenStudy (anonymous):

yes that's it "/

OpenStudy (anonymous):

\[\sqrt{5y+1}=7\]

OpenStudy (saifoo.khan):

5y + 1 = 49

OpenStudy (anonymous):

this isn't the simplified version...correct?

OpenStudy (anonymous):

y=48/5?